media%2F267%2F267dc669-80de-4744-a28f-57
media%2Fb02%2Fb0281149-f84a-4957-a992-14
media%2F955%2F9553d901-7998-44b9-b968-16
media%2Fd53%2Fd53f58ce-9cc2-4bf2-bc43-6a
5. 20 points The following summaries from the income statements and balance sheets of Kouris Company and Brittania, Inc. are presented below. Round to two decimal places. Show your math For both companies for 2016, compute the: Brittania Kouris Current Ratio Acid Test Ratio Accounts Receivable Turnover Inventory Tunover Days' Sales in Inventory Days Sales Uncollected Profit Margin Ratio Return on Total Assets Return on Common Stockholders'Equity Which company do you consider to be the better short-term credit risk? Which company do you consider to have better profitability ratios?
Answer #1
Kouris Brittania
(a) Current ratio:-
a. Current Assets $   4,679.9 $   80.7
b. Current Liabilities $   2,015.2 $   16.3
Ratio (a/b) $     2.32 times 4.95 times
(b) Acid Test Ratio:-
a. Quick Assets
Total current assets- Inventory 4679.9-1514.9 = $   3,165.0 80.7-27.2 = $   53.5
b. Current Liabilities $   2,015.2 $   16.3
Ratio (a/b) 1.57 times 3.28 times
(c) Accounts Receivable turnover :-
a. Net Credit Sales $ 10,697.0 $ 133.5
b. Average Accounts Receivable
opening + closing 1804.1+2101.1 = $   1,952.6 14.7+15.5 = $   15.1
2 2 2
Ratio (a/b) 5.48 times 8.84 times
(d) Inventory turnover:-
a. Cost of goods sold $   6,313.6 $   87.3
b. Average Inventory
opening + closing 1373.8+1514.9 = $   1,444.4 28.4+27.2 = $   27.8
2 2 2
Ratio (a/b) 4.37 times 3.14 times
( e) Days' Sales in Inventory:-
a. Number of Days in year 365 365
b. Inventory turnover ratio 4.37 times 3.14 times
Ratio (a/b) 83.50 days 116.23 days
(f) Days Sales uncollected:-
a. Number of Days in year 365 365
b. Accounts Receivable turnover ratio 5.48 times 8.84 times
Ratio (a/b) 66.63 days 41.28 days
(g) Profit Margin Ratio:-
a. Net Income before effect of accounting change $      740.1 $     5.2
b. Net Sales $ 10,697.0 $ 133.5
Ratio (a/b) 6.92% 3.90%
(h) Return on Total Assets Ratio:-  
a. Net Income before effect of accounting change $      740.1 $   5.20
b. Interest Expense (1-tax rate) 42.9 * (1- (382.9/1123)) $        28.3 0.1* (1- (3.9/9.1)) $   0.06
c Net income + Interest Expense (1-tax rate) $      768.4 $   5.26
b. Average Total assets
opening + closing 6440+6713.9 = $   6,577.0 78+87.5 = $      83
2 2 2
Ratio (a/b *100) 11.68% 6.35%
(h) Return on common stockholders' equity:-
a. Net Income $         474 $     5.2
b. Common Stockholders' Equity
opening + closing 3839+3990.7 = $      3,915 63.2+68.7 = $   66.0
2 2 2
Ratio (a/b *100) 12.11% 7.88%
Brittania Company - better short term credit risk
Kouris Company - better profitability raios

Feel free to ask any clarification, if required. Please provide feedback by thumbs up, if satisfied. It will be highly appreciated. Thank you.